OpenStudy (anonymous):

Linda invests $13,000 for one year. Part is invested at 4%, another part at 5%, and the rest at 7%. The total income from all 3 investments is $705. The combined income from the 4% and 5% investments is $75 more than the income from the 7% investment. Find the amount invested at each rate (can somebody please help me solve this problem, i am completely lost)

OpenStudy (anonymous):

cant be too bad

OpenStudy (anonymous):

let amount invested at 4% be x ammount at 5% be y amount at 7% be z

OpenStudy (anonymous):

we know that the total amount invested is 13000 so x+y+z = 13000 ( one eqn )

OpenStudy (anonymous):

now, assuming simple interest ( which I think is what they want )

OpenStudy (anonymous):

I = Prn

OpenStudy (anonymous):

income from $x = x (0.04)(1) =0.04x

OpenStudy (anonymous):

x*1.04+y*1.05=z*1.07 With only one year it has to be simple interest.

OpenStudy (anonymous):

so ^ thats another eqn there

OpenStudy (anonymous):

and last condition gives another eqn 0.04x +0.05y = 75 + 1.07z

OpenStudy (anonymous):

lol that guys eqn was wrong

OpenStudy (anonymous):

yeah it was, sorry, fiddling about now

OpenStudy (anonymous):

should have been 0.04x + 0.05y + 0.07z = 705

OpenStudy (anonymous):

so x+y+z = 13000 0.04x + 0.05y + 0.07z = 705 0.04x +0.05y = 75 + 1.07z

OpenStudy (anonymous):

there eqns, three unknowns you can solve, im not doing boring pointless albegra

OpenStudy (anonymous):

algebra

OpenStudy (anonymous):

no EDIT : x+y+z = 13000 0.04x + 0.05y + 0.07z = 705 0.04x +0.05y = 75 + 0.07z made a little mistake when copying the second eqn

OpenStudy (anonymous):

they should have it so you can delete posts z :| also, if you were to want to solve these eqns, the best first step would be to subtract eqns 2 and 3

OpenStudy (anonymous):

yeah just word problems arent my thing :(